OpenStudy (anonymous):

|3x+12|=18 The solution to this equation is 2. Answer a.True b.False

OpenStudy (mertsj):

It seems like there should be another answer as well since there are 2 numbers that are exactly 18 units from 0. What do you think?

OpenStudy (anonymous):

false?

OpenStudy (mertsj):

I think so.

OpenStudy (anonymous):

\[|3x+12|=18\] \[3x+12=\pm18\] The equations splits off here in two parts, 3x+12=18 and 3x+12=-18 This means you MUST have two solutions unless you can eliminate one based on conditions (i.e.: the answer has to be positive) They didn't give you conditions, so you have more than one answer. False. :-)

OpenStudy (anonymous):

By the way, can you find what the other answer is for that equation? ;-)